Department E had 4,000 units in Work in Process that were 40% completed at the beginning of the period at a cost of $12,500. 14,000 units of direct materials were added during the period at a cost of $28,700. 15,000 units were completed during the period, and 3,000 units were 75% completed at the end of the period. All materials are added at the beginning of the process. Direct labor was $32,450 and factory overhead was $18,710. The number of equivalent units of production for the period for materials if the average cost method is used to cost inventories was a. 18,000 b. 17,250 c. 15,650 d. 17,700
Answer #1

Number of equivalent units of production

(a) Number of completed units during the period = 15000 units ( 100 % completion)

(b) Number of partly completed units during the period = 3000 Units ( 75 % of completion)

(c) Equivalent number of partly completed units = 3000 X 75% = 2250 units

(d) Total number of equivalent units of production = (a) + (c) = ( 15000 + 2250 ) = 17,250 units

Total no. of equivalent units of production for the period is 17250 units

No. of units introduced No. of units produced
Units in W I P ( partially completed) 4000 (40% completed) 3000 ( 75% completed)
Finished units (Fully completed) 14000 (100% Completed) 15000 ( 100% Completed)
Total( General units) 18000 units 18000 units
Total equivalent production units

15000 + (3000*75%)

= 15000 + 2250

= (b) 17,250 units